# Stability of time-periodic solutions to the Navier-Stokes-Fourier system

### v1

@@ Abstract
We prove the existence and stability of a time-periodic solution to the Navier-Stokes-Fourier system in the three-dimensional whole space when a time-periodic external force is sufficiently small. The time decay estimate of the perturbation around the time-periodic solution is derived under the assumption that an initial perturbation is small and belongs to $L^{p}$ for some $1\leq p\leq 2$.
@@ 1 Introduction
$\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0,\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f,\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho}.\end{aligned}\right$. (1)
Here, $f$ is the external body force, $P$ is the pressure, $\mu$ and $\mu^{\prime}$ are the viscous coefficients, $\kappa$ is the heat conductivity coefficient, $c_{V}$ is the specific heat at constant volume and $\Psi=\Psi(v)$ is the dissipation function given by
@@
The aim of this paper is twofold. The first is to investigate the existence of time-periodic solutions to the Navier-Stokes-Fourier system in $\mathbb{R}^{3}$ subject to a given small time-periodic external force. The second is to examine the stability of the time-periodic solutions with respect to external forces and initial perturbations, respectively. Let $f$ be a time-periodic force with time-period $T>0$, that is,
@@
We consider the time-periodic problem (2) around the constant state
@@
The time-periodic problem of the Navier-Stokes-Fourier system was considered by Ma, Ukai and Yang [MR2595722] in $\mathbb{R}^{N}$, $N\geq 5$. They proved the existence and stability result for the time-periodic problem by effectively applying the energy method in the inhomogeneous Sobolev space of $L^{2}$ type, but the dimensional assumption $N\geq 5$ is needed to control the $L^{2}$ norm of the time-periodic solution. Indeed, since the diffusion terms
are classified as elliptic operators, the spatial decay orders of the velocity $v_{T}$ and the temperature $\theta_{T}-\theta_{\infty}$ are expected at most $O(1/|x|^{N-2})$ as $|x|\to\infty$, respectively, by comparing the spatial decay order of the Newtonian potential. See also the asymptotic expansion result for the incompressible Navier-Stokes equation in [MR1794517]. If the dimension $N=3$, then the expected spatial decay order is $O(1/|x|)$, which does not belong to $L^{2}$, so that this slow spatial decay suggests a need for a new function framework to construct a time-periodic solution in dimension $N=3$. In order to establish a function space setting that admits slow spatial decay, such as $1/|x|$, we introduce the homogeneous Besov spaces as follows. Let $s\in\mathbb{R}$ and $1\leq r\leq\infty$. Choose $\phi\in C^{\infty}$ supported in the annulus $\mathcal{C}=\{\xi\in\mathbb{R}^{3}\mid 3/4\leq|\xi|\leq 8/3\}$ such that
$\sum_{j\in\mathbb{Z}}\phi(2^{-j}\xi)=1\ \ \ \mathrm{for}\ \ \ \xi\neq 0.$
$\dot{\Delta}_{j}u\equiv\mathcal{F}^{-1}[\phi(2^{-j}\cdot)\hat{u}$].
with norm
where $\mathcal{S}^{\prime}$ is the set of all Schwartz distributions and $\mathcal{P}$ is the polynomial ring over $\mathbb{R}$. We shall employ a function space to construct the time-periodic solution as the space $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\in\mathbb{Z}_{\geq 5}$, which spaces allow for slow decay, such as $1/|x|$, and have good compatibility with the energy method, since the space $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$ is $L^{2}$-based space. Here, we mention that the space $\dot{B}^{1/2}_{2,1}\cap\dot{H}^{3}$ was employed for the perturbation analysis of stationary solutions to the isentropic compressible Navier-Stokes equations in [MR4803477].
Let us state our main theorem, which shows the existence and stability of the time-periodic problem (2).
@@
then the following assertions hold.
@@
The time-periodic problem for the incompressible Navier-Stokes equation has been extensively studied; we refer here to Maremonti [MR1107016], Kozono-Nakao [MR1373173], Yamazaki [MR1777114], Galdi-Silvestre [MR2603768], Kobayashi [MR2589952] and Kučera [MR2610568] among many others. In particular, Yamazaki [MR1777114] derived time-decay estimates for perturbations around a time-periodic solution to the incompressible Navier-Stokes equation in unbounded domains such as the whole space, the half-space, or exterior domains of dimension $N\geq 3$. For the isentropic compressible Navier-Stokes equation, the existence and stability of the time-periodic solution were studied by Valli [MR753158] in a 3D bounded domain, and by Kagei-Tsuda [MR3274764], Tsuda [MR3437858] in $\mathbb{R}^{N}$, $N\geq 3$. The existence and stability of the time-periodic solution to the Navier-Stokes-Fourier system was studied by Ma, Ukai and Yang [MR2595722] in $\mathbb{R}^{N}$, $N\geq 5$. We also mention that Feireisl et al. [MR2917121] show the existence of at least one weak time periodic solution of the Navier-Stokes-Fourier system in a 3D bounded domain for the specific pressure. Since the spatial decay order of the time-periodic solution is expected to be at most O(1/|x|){O}(1/|x|), the convection term $v\cdot\nabla v$ appearing in the Navier-Stokes equations is expected to exhibit the same spatial decay order, ${O}(1/|x|^{3})$, as the viscous term $\mu\Delta v$. Thus, we need to treat the convection terms carefully to drive the existence of a time-periodic solution even if the given external force is small enough. For this reason, Yamazaki and Tsuda transform the convection term $v\cdot\nabla v$ to the divergence forms $\mathrm{div}(v\otimes v)$ and $\mathrm{div}(\rho v\otimes v)$, respectively, by using the mass conservation low, and then establish a priori estimate for the solutions. In the Navier-Stokes-Fourier system case, the additional convection term $v\cdot\nabla\theta$ and the non-divergence form term $(\theta\partial_{\theta}P/c_{V}\rho)\mathrm{div}v$ appear in the equation of the temperature in (1). In Section 3 below, to establish a priori estimate in the homogeneous Besov spaces, we shall introduce the modified energy functional
@@
where $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are the higher order terms. Except for the external force terms and the higher-order terms, all nonlinear terms in (5) are written in divergence forms or potential forms. This transform is employed to establish a priori estimates for the initial value problem (1) in the proofs of Propositions 3.5 and 3.6 below. The proof of the time-decay estimate (((ii))) is inspired by the strategy employed in [MR4803477]. The key step of the proof is to derive the time-weighted estimate for the Duhamel integral:
@@
$\hskip 50.0pt\lesssim\sup_{s_{1}\leq s\leq\frac{3}{2}-\epsilon}\ \sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\|\Phi(t)\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$,
where $\epsilon>0$, $3/2\leq s_{0}\leq 1/2$, $s_{1}=\max\{s_{0},0\}$, $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$ and $e^{tA}$ is the linearized semigroup of (5) around the constant state which defined in (12) below. A similar analysis is found in Yamazaki [MR1777114].
The outline of this paper is as follows: In Section 2, we present some lemmas related to the homogeneous Besov space, which will be frequently used in the subsequent sections. Section 3 is devoted to the proof of Theorem 1.1. The proof of Theorem 1.1 is established by combining the local existence theorem with a priori estimates for the solutions to the initial value problem of (1). A priori estimate and the local existence result for the initial value problem are proved in Subsection 3.1 and Subsection 3.2, respectively. We show the uniform in time and the time-weighted a priori estimate for the solution in Proposition 3.1 and Proposition 3.2 below, respectively. A priori estimates are performed for the homogeneous Besov norms and the homogeneous Sobolev norms in a parallel way. The proof of the estimate in the homogeneous Besov spaces is carried out by using the modified energy functional introduced in (1). Here, the time-space norm estimate in (18) below for the linearized semigroup of (5) around the constant state $(\rho_{\infty},0,\theta_{\infty})$ plays a crucial role in deriving the homogeneous Besov norm estimates. (Cf. [MR1777114], [MR4803477].) The proof of the estimates in the homogeneous Sobolev norms is established by the energy method. The proof of the local existence is carried out by constructing the approximating sequence of the solution. The approximating sequence is constructed as the solution of the transport equation for the density and as the solution of the quasilinear parabolic equation for the velocity and the temperature, since the density causes the quasilinear effect in the equations of the velocity and the temperature in (1). By virtue of employment in this construction, we obtain the local existence result without the smallness assumption for initial data, unlike the local existence result in [MR4803477]. Finally, the proof of Theorem 1.2 is given in Subsection 3.3.
@@
For $I=[0,T_{0})$ with $0<T_{0}\leq\infty$, we write $L^{p}_{T_{0}}(Z)$ and $C_{T_{0}}(Z)$ instead of $L^{p}(I;Z)$ and $C(I;Z)$, respectively. In particular, if $T_{0}=\infty$, we simply write $L^{p}(Z)$ and $C(Z)$ by omitting the subscript.
@@ 2 Preliminary
In this section, we present several lemmas related to the homogeneous Besov space, which are frequently used in this paper. The following basic properties for the homogeneous Besov space are proved in [MR2768550, Lemma 2.1, Proposition 2.22, Theorem 2.25 and Proposition 2.29].
@@
The following embedding property holds. (Cf. [MR2768550, Proposition 2.20, Propositon 2.39 and Theorem 2.40].)
@@
The following commutator estimate is the special case of [MR2768550, Lemma 2.100 and Remark 2.102].
@@
We state the bilinear estimate in the homogeneous Besov space. The following lemma is a direct consequence of [MR2768550, Theorem 2.47 and Theorem 2.52].
@@
We use the following lemmas regarding the composition of functions. (Cf. [MR2768550, Theorem 2.61].)

### v2

# Stability analysis of time-periodic solutions to the Navier-Stokes-Fourier system in 3D whole space

@@ Abstract
This paper concerns the large-time behavior of perturbations around a time-periodic solution to the Navier-Stokes-Fourier system in the three-dimensional whole space. The time-periodic solution exists when a given external force is small enough. We derive the time-decay estimate of the perturbation under the assumption that an initial perturbation is sufficiently small. The time-space integral estimate for the linearized semigroup around the constant state in the Besov spaces is effectively applied in the proof.
@@ 1 Introduction
$\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho}\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3}.\end{aligned}\right$. (1)
Here, $f=f(t,x)$ is the external body force, $P$ is the pressure, $\mu$ and $\mu^{\prime}$ are the viscous coefficients, $\kappa$ is the heat conductivity coefficient, $c_{V}$ is the specific heat at constant volume and $\Psi=\Psi(v)$ is the dissipation function given by
@@
Let $f$ be a time-periodic force with time-period $T>0$, that is,
@@
In this paper, we consider the time-periodic problem (2) around the constant state
@@
The time-periodic problem has been extensively studied for the incompressible Navier-Stokes equation; we refer here to Maremonti [12], Kozono-Nakao [9], Yamazaki [17], Galdi-Silvestre [5], Kobayashi [8] and Kučera [10] among many others. In particular, Yamazaki [17] derived time-decay estimates for perturbations around a uniformly bounded in time solution, including a time-periodic solution, to the incompressible Navier-Stokes equation in unbounded domains such as the whole space, the half-space, or exterior domains of dimension $n\geq 3$ under the assumption of smallness for both the external force and the initial perturbation. For the isentropic compressible Navier-Stokes equation, the existence and stability of the time-periodic solution were studied by Valli [16] in a 3D bounded domain, and by Kagei-Tsuda [6], Tsuda [15] in $\mathbb{R}^{n}$, $n\geq 3$. For the non-isentropic case, Ma, Ukai and Yang [11] proved the existence and stability result for the compressible Navier-Stokes system in $\mathbb{R}^{n}$, $n\geq 5$ and Tsuda proved the existence and stability result for the compressible Navier-Stokes-Korteweg system in $\mathbb{R}^{3}$. We also mention that Feireisl et al. [4] show the existence of at least one weak time periodic solution to the Navier-Stokes-Fourier system in a 3D bounded domain for large data. The time decay estimate of the perturbation around a time-periodic solution was derived by Ma, Ukai and Yang [11] for the Navier-Stokes-Fourier system under the smallness assumptions. They prove the decay estimate
$\|(\rho-\rho_{T},v-v_{T},\theta-\theta_{T})(t)\|_{H^{N-1}}$
$\hskip 50.0pt\lesssim(1+t)^{-\frac{n}{4}}\|(\rho_{0}-\rho_{T}(0),v_{0}-v_{T}(0),\theta_{0}-\theta_{T}(0))\|_{L^{1}\cap H^{N-1}},$ (3)
where $N\in\mathbb{Z}_{\geq n+2}$ and $H^{N-1}$ is the inhomogeneous Sobolev space defined in the end of this section. Here, $(\rho,v,\theta)$ is a global solution to (1) in $\mathbb{R}^{n}$, $n\geq 5$ with initial data $(\rho_{0},v_{0},\theta_{0})$ and $(\rho_{T},v_{T},\theta_{T})$ is the corresponding time-periodic solution. The decay rate in (1) corresponds to that of the heat semigroup. However, the proof of decay estimate in [11] relies on the dimensional assumption $n\geq 5$. In this paper, we aim to establish the decay estimates for perturbations around a time-periodic solution to the Navier-Stokes-Fourier system in the physically important three-dimensional case.
Let us state our main theorem, which derives the decay estimate for the perturbation.
@@
then the following assertions hold. Here, the space $\dot{B}_{2,\infty}^{\frac{1}{2}}$ is the Besov space defined in Section $2$ below.
@@
We note that the decay rate obtained in (((ii))) coincides with that of the heat semigroup. The employment of the Besov space is motivated by the need to control the slow spatial decay of the time-periodic solution. Indeed, since the diffusion terms
are classified as elliptic operators, the spatial decay orders of the velocity $v_{T}$ and the temperature $\theta_{T}-\theta_{\infty}$ are expected at most $O(1/|x|)$ as $|x|\to\infty$, respectively, by comparing the spatial decay order of the Newtonian potential in $\mathbb{R}^{3}$. This slow decay suggests that the time-periodic solution does not belong to $L^{2}(\mathbb{R}^{3})$. Previously, Ma, Ukai and Yang [11] derived the decay estimate (1) in dimension $n\geq 5$ by combining the analysis using the linearized semigroup around the constant state $(\rho_{\infty},0,\theta_{\infty})$ and the energy estimates in $H^{N-1}$, $N\in\mathbb{Z}_{\geq n+2}$. In their work, the dimensional assumption $n\geq 5$ is used to control the $L^{2}$ norm of the time-periodic solution. To establish the decay estimate of the perturbation in dimension $n=3$, we develop the framework of perturbation analysis in the function spaces $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\geq\mathbb{Z}_{\geq 5}$, which spaces allow for slow decay, such as $1/|x|$. In addition, the spaces $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\geq\mathbb{Z}_{\geq 5}$ have good compatibility with the energy method developed by Matsumura and Nishida [14] for constructing strong solutions to the compressible Navier-Stokes-Fourier system, since the spaces are $L^{2}$-based space. Here, we mention that the space $\dot{B}^{1/2}_{2,1}\cap\dot{H}^{3}$ was used for the perturbation analysis of stationary solutions to the isentropic compressible Navier-Stokes equations in [3].
The difficulty in proving Theorem 1.1 arises from the convection terms in the Navier-Stokes-Fourier system. Since the spatial decay order of the time-periodic solution is expected to be at most O(1/|x|){O}(1/|x|), the convection term $v\cdot\nabla v$ appearing in the Navier-Stokes equations is expected to exhibit the same spatial decay order, ${O}(1/|x|^{3})$, as the viscous term $\mu\Delta v$. Thus, we need to treat the convection terms carefully to prove the existence of a time-periodic solution and to derive the decay estimate for its perturbations even if the given external force is small enough. For this reason, Yamazaki and Tsuda transform the convection term $v\cdot\nabla v$ to the divergence forms $\mathrm{div}(v\otimes v)$ and $\mathrm{div}(\rho v\otimes v)$, respectively, by using the mass conservation low, and then establish a priori bound for the solutions to the Navier-Stokes equations. However, in the Navier-Stokes-Fourier system case, the additional convection term $v\cdot\nabla\theta$ and the non-divergence form term $(\theta\partial_{\theta}P/c_{V}\rho)\mathrm{div}v$ appear in the equation of the temperature in (1). In Section 3 below, to establish a priori estimate in the homogeneous Besov spaces, we shall introduce the modified energy functional
@@
where $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are the higher order terms. Except for the external force terms and the higher-order terms, all nonlinear terms in (6) are written in divergence forms or potential forms. This transform is used to establish a priori estimates for the initial value problem (1) in the proofs of Propositions 3.5 and 3.6 below. The proof of the time-decay estimate (((ii))) is inspired by the strategy used in [3]. The proof will be carried out by writing the perturbation $U=(\sqrt{p_{1}}\sigma,\ m,\ \sqrt{c_{V}/\theta_{\infty}}\mathcal{E})$ as the Duhamel form:
$U(t)=e^{tA}U(0)+\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau$
where $t>0$,
$A=\left[\begin{matrix}0&-\sqrt{p_{1}}\mathrm{div}&0\\ -\sqrt{p_{1}}\nabla&\frac{\mu}{\rho_{\infty}}\Delta+\frac{\mu+\mu^{\prime}}{\rho_{\infty}}\nabla\mathrm{div}&-\frac{p_{2}}{\rho_{\infty}}\sqrt{\frac{\theta_{\infty}}{c_{V}}}\nabla\\ 0&-\frac{p_{2}}{\rho_{\infty}}\sqrt{\frac{\theta_{\infty}}{c_{V}}}\mathrm{div}&\frac{\kappa}{c\rho_{\infty}}\Delta\end{matrix}\right],$
and $\Phi$ denotes the remainder terms that contain the linear terms such as $\mathrm{div}(v_{T}\otimes m)$. The key step of the proof is to derive the time-weighted estimate for the Duhamel integral:
@@
$\hskip 50.0pt\lesssim\sup_{s_{1}\leq\tilde{s}\leq\frac{3}{2}-\epsilon}\ \sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{\tilde{s}-s_{0}}{2}}\|\Phi(t)\|_{\dot{B}^{\tilde{s}-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$,
where $\epsilon>0$, $3/2\leq s_{0}\leq 1/2$, $s_{1}=\max\{s_{0},0\}$, $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$. By virtue of rewriting the convection term as the divergence form, the bilinear estimate in Lemma 2.5 below allow us to prove the estimates such as
$\|\mathrm{div}(v_{T}\otimes m)\|_{\dot{B}^{\tilde{s}-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}\lesssim\|v_{T}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{3}}\|m\|_{\dot{B}^{\tilde{s}}_{2,\infty}\cap\dot{H}^{3}},\quad s_{1}\leq\tilde{s}\leq 3/2-\epsilon,$
so that if the time-periodic solution is sufficiently small and the high-frequency norm $\dot{H}^{3}$ of the perturbation is well controlled, then the linear operator $\mathrm{div}(v_{T}\otimes\cdot)$ can be treated as a simple perturbation of the operator $A$. The time-weighted estimate in the high-frequency norm is proved by using the energy method developed by Matsumura and Nishida [14] in Proposition 3.8 below.
The outline of this paper is as follows: In Section 2, we present some lemmas related to the homogeneous Besov space, which will be frequently used in the subsequent sections. Section 3 is devoted to the proof of Theorem 1.1. The proof of Theorem 1.1 is established by combining the local existence theorem with a priori estimates for the solutions to the initial value problem of (1). A priori estimate and the local existence result for the initial value problem are proved in Subsection 3.1 and Subsection 3.2, respectively. We show the uniform in time and the time-weighted a priori estimate for the solution in Proposition 3.1 and Proposition 3.2 below, respectively. A priori estimates are performed for the homogeneous Besov norms and the homogeneous Sobolev norms in a parallel way. The proof of the estimate in the homogeneous Besov spaces is carried out by using the modified energy functional introduced in (1). Here, the time-space norm estimate in (19) below for the linearized semigroup of (6) around the constant state $(\rho_{\infty},0,\theta_{\infty})$ plays a crucial role in deriving the homogeneous Besov norm estimates. (Cf. [17], [3].) The proof of the estimates in the homogeneous Sobolev norms is established by the energy method. The proof of the local existence is carried out by constructing the approximating sequence of the solution. The approximating sequence is constructed as the solution of the transport equation for the density and as the solution of the quasilinear parabolic equation for the velocity and the temperature, since the density causes the quasilinear effect in the equations of the velocity and the temperature in (1). By virtue of employment in this construction, we obtain the local existence result without the smallness assumption for initial data, unlike the local existence result in [3]. Finally, the proof of Theorem 1.1 is given in Subsection 3.3.
@@
For $I=[0,T_{0})$ with $0<T_{0}\leq\infty$, we write $L^{p}_{T_{0}}(Z)$ and $C_{T_{0}}(Z)$ instead of $L^{p}(I;Z)$ and $C(I;Z)$, respectively. In particular, if $T_{0}=\infty$, we simply write $L^{p}(Z)$ and $C(Z)$ by omitting the subscript. For any $s\in\mathbb{R}$, $H^{s}=H^{s}(\mathbb{R}^{3})$ denotes the inhomogeneous Sobolev space by
$H^{s}=\{u\in\mathcal{S}^{\prime}\mid(1+|\cdot|^{2})^{\frac{s}{2}}\hat{u}\in L^{2}\}$
with norm
$\|u\|_{H^{s}}=\|(1+|\cdot|^{2})^{\frac{s}{2}}\hat{u}\|_{L^{2}}.$
@@ 2 Preliminary
In this section, we present the definition and several lemmas for the homogeneous Besov space, which are frequently used in this paper. The following basic properties for the homogeneous Besov space are proved in [1, Lemma 2.1, Proposition 2.22, Theorem 2.25 and Proposition 2.29].
Definition 2.1.
Let $s\in\mathbb{R}$ and $1\leq r\leq\infty$. Choose $\phi\in C^{\infty}$ supported in the annulus $\mathcal{C}=\{\xi\in\mathbb{R}^{3}\mid 3/4\leq|\xi|\leq 8/3\}$ such that
$\sum_{j\in\mathbb{Z}}\phi(2^{-j}\xi)=1\ \ \ \mathrm{for}\ \ \ \xi\neq 0.$
$\dot{\Delta}_{j}u\equiv\mathcal{F}^{-1}[\phi(2^{-j}\cdot)\hat{u}$].
with norm
where $\mathcal{P}$ is the polynomial ring over $\mathbb{R}$.
The following lemma shows the basic properties of the Besov space.
@@
The following embedding property holds. (Cf. [1, Proposition 2.20, Propositon 2.39 and Theorem 2.40].)
@@
The following commutator estimate is the special case of [1, Lemma 2.100 and Remark 2.102].
@@
We state the bilinear estimate in the homogeneous Besov space. The following lemma is a direct consequence of [1, Theorem 2.47 and Theorem 2.52].
@@
We use the following lemmas regarding the composition of functions. (Cf. [1, Theorem 2.61].)